OpenStudy (littlet):

If Sin x = 0.63 what is the value of x? explain the steps using a scientific calculator.

OpenStudy (littlet):

@phi

OpenStudy (phi):

didn't you do this the idea is to use inverse sin (on your calculator do you have a key \( \sin^{-1}\) ?

OpenStudy (littlet):

Yes i do

OpenStudy (phi):

the inverse sin undoes the sin you start with [ \sin x = 0.63\] you "apply" the inverse sign to both sides by writing \[ \sin^{-1} (\sin x) = \sin^{-1} 0.63 \] the left side is just x (because inverse sin of sin undoes it) in other words, we have \[ x= \sin^{-1} 0.63\] so you do inverse sin of 0.63 on your calculator what do you get ?

OpenStudy (littlet):

i got 39.050 is that right?

OpenStudy (phi):

yes you can check by doing sin(39.05) and that should give back 0.63

OpenStudy (littlet):

Yes that worked. Thank you again!! have a great night!